Two identical spheres of radius 30mm and mass 2kg on the surface of the earth are placed in contact. Find the gravitational attraction between them.
1
Expert's answer
2020-09-22T15:35:15-0400

The  gravitational attraction between two spheres is given by the Newton's gravitational law:


"F = G\dfrac{m_1m_2}{r^2}"

where "m_1 = m_2 = 2kg" are masses of two identical spheres, "r = 2\cdot 30mm = 0.06m" is the distance between their centers and "G = 6.67\times10^{-11} m^3kg^{-1}s^{-2}" is the gravitational constant. Thus, obtain:

"F = G\dfrac{m_1m_2}{r^2} = 6.67\times10^{-11} \cdot \dfrac{2\cdot 2}{0.06^2} \approx 7.41\times 10^{-8}N"

Answer. "7.41\times 10^{-8}N".